OpenStudy (anonymous):

The capacity, c, of a set of similar containers is directly proportional to the cube of their heights,h. If a container of height 20cm holds 2 litres, find a) the capacity of a container which is 14cm high b) the height of a container which has a capacity of 3456cm^3

OpenStudy (texaschic101):

20/2 = 14/x --- 20 cm to 2 liters = 14 cm to x liters cross multiply 20x = 28 x = 28/20 x = 1 8/20 = 1 2/5 so 14 cm holds 1 and 2/5 liters sorry...don't know the second one